OpenStudy (rock_mit182):

HELp I need to factor out : \[x ^{4} + 4x^{3} + 12x^{2} + 16x+16\] I

OpenStudy (rock_mit182):

I got the answer but actually i do not know how to get there

OpenStudy (rock_mit182):

\[x ^{4} + 4x^{3} + 12x^{2} + 16x+16\]

OpenStudy (rock_mit182):

@quickstudent @mathmate @pooja195

OpenStudy (rock_mit182):

@IrishBoy123

OpenStudy (rock_mit182):

@inkyvoyd

OpenStudy (quickstudent):

Oh sorry, I don't exactly remember how to do this

OpenStudy (rock_mit182):

:/

OpenStudy (rock_mit182):

@Conqueror

OpenStudy (rock_mit182):

any ideas ?

OpenStudy (rock_mit182):

@whpalmer4

OpenStudy (sshayer):

\[x^4+4x^3+12x^2+16x+16=x^2(x^2+4x+12+\frac{ 16 }{ x }+\frac{ 16 }{ x^2 })\] \[=x^2\left\{ \left( x2+\frac{ 16 }{ x^2 } \right)+\left( 4x+\frac{ 16 }{ x } \right)+12 \right\} ...(1)\] \[put~x+\frac{ 4 }{ x }=t\] squaring \[x^2+\frac{ 16 }{ x^2 }+8=t^2,x^2+\frac{ 16 }{ x^2 }=t^2-8\] \[=x^2[(t^2-8)+4t+12]=x^2[t^2+4t+4]=x^2(t+2)^2\] \[=x^2(x+\frac{ 4 }{ x }+2)^2=[x(x+\frac{ 4 }{ x }+2)]^2=(x^2+2x+4)^2\]

OpenStudy (rock_mit182):

DUDE unbelievable!

OpenStudy (rock_mit182):

really, thanks bro

OpenStudy (sshayer):

yw